OpenStudy (wesdg1978):

Express the given function as a composite of functions f and g such that y=f(g(x)). y=|5x+8| & y=2/sqrt(6x+9)

OpenStudy (zzr0ck3r):

1) f(x) = |x| and g(x) 5x+8 2) f(x)=2/x and g(x) = sqrt(6x+9)

OpenStudy (zzr0ck3r):

f(x)=|5x+8|, g(x) = x would also work but its trivial

OpenStudy (wesdg1978):

Can you show me how you got that?

OpenStudy (zzr0ck3r):

f(x) = |x| and g(x) = 5x+8 so f(g(x)) = f(5x+8) = |5x+8|